# Shaocheng Liu and Qi Chen

### v2

@@ IV-B Entropy functions on the faces extended from $2$-dimensional faces of $\Gamma_{3}$
By restriction, the above $3$ faces of $\Gamma_{4}$ are mapped to a subset of $\Gamma_{3}$ on $\{2,3,4\}$. The face $(\hat{U}_{2,5}^{1},U_{1,1}^{1})$ of $\Gamma_{4}$ is mapped to the extreme ray $U_{2,3}$ of $\Gamma_{3}$ on $\{2,3,4\}$, which implies that $a$ can only take $\log k$ for integer $k>0$. For the faces $(\hat{U}_{2,5}^{1},U_{1,1}^{2})$ and $(\hat{U}_{2,5}^{1},U_{1,2}^{12})$, by restriction, they are both mapped to the face $(U_{2,3},U^{1}_{1,1})$ of $\Gamma_{3}$ on $\{2,3,4\}$, which implies that $a$ can only take $\log k$ for positive integer $k$ as well.
For the “if ” part, as the whole ray $U_{1,1}^{i}$ and $U_{1,2}^{12}$ are entropic, by Lemma 2 and Theorem II-C.4, all $a=\log k$ are entropic. ∎
@@
Let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are the rank functions of the minimal integer polymatroids on the two extreme rays of the face, respectively. Restricting $\mathbf{h}$ on $\{2,3,4\}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}_{1}+b\mathbf{r}^{\prime}_{2}$, where $\mathbf{r}^{\prime}_{i},i=1,2$ are the restriction of $\mathbf{r}_{i}$ on $\{2,3,4\}$. By restriction, The face $(\hat{U}_{2,5}^{1},U_{1,3}^{123})$ of $\Gamma_{4}$ is mapped to the face $(U_{2,3},U^{12}_{1,2})$ of $\Gamma_{3}$ on $\{2,3,4\}$, Thus, $\mathbf{h}^{\prime}\in(U_{2,3},U^{12}_{1,2})$ and $a+b\geq\log\lceil 2^{a}\rceil$ by [9, Theorem 1].
For the “if” part , we only need to show that $\mathbf{h}$ with $a+b=\log k$ is entropic for each positive integer $k$, then the theorem is implied by Lemma 2 and the fact that $b\mathbf{r}_{2}$ is entropic for all $b\geq 0$. Let $X_{2}$ be uniformly distributed on $\{0,1,\dots,k-1\}$. Let $X_{2}$ and $X_{4}$ be independent, and $X_{4}$ be distributed on $\{0,1,\dots,k-1\}$ such that $H(X_{4})=a$. Let $X_{1}=kX_{2}+X_{4}$, $X_{3}=X_{2}+X_{4}\mod k$. It can be checked that the entropy function of such constructed $(X_{i},i\in N_{4})$ is $(a,\log k-a)$. ∎
@@
For $F=(\hat{U}_{3,5}^{1},U_{1,1}^{1})$, $(\hat{U}_{3,5}^{1},U_{1,1}^{2})$, or $(\hat{U}_{3,5}^{1},U_{1,2}^{12})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log k$ for some integer $k>0$.
@@
For $x_{134}$ with $p(x_{134})>0$, above information equalties imply that the probability mass function satisfies
$p(x_{1}.x_{3},x_{4})$ $=p(x_{1},x_{3})=p(x_{1}.x_{4})$
@@
which implies $p(x_{3})=p(x_{4})$. Since $X_{3}$ and $X_{4}$ are independent, by Lemma 1, they are uniformly distributed on $\mathcal{X}_{3}$ and $\mathcal{X}_{4}$, respectively, and $H(X_{3})=H(X_{4})=\log k$, where $k=|\mathcal{X}_{3}|=|\mathcal{X}_{4}|$. As $(X_{i},i\in\{1,3,4\})$ is its characterizing random vector, we have
@@
Thus $a$ can only take $\log k$.
For the ”if” part, it can be proven by Lemma 2 and the fact that $a=\log{k}$ in $\hat{U}_{3,5}^{1}$ and whole ray $U^{1}_{1,1}$, $U^{2}_{1,1}$ and $U_{1,2}^{12}$ are entropic. ∎
@@ IV-C Entropy functions on the faces involving $U_{2,3}^{123}$
For $F=(U_{2,3}^{123},U_{1,3}^{124})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a$= $\log{k}$ for integer $k>0$.
$a$$b$Olog2log3log4
@@
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),i=3,j\in N_{4}\setminus\{3\},$
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),k\neq 3,\{i,j,k\}\neq\{1,2,3\},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),K\subseteq\{1,2,3\},|K|=2,\{i,j\}=N_{4}\setminus K.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
@@
Consider the bipartite graph $G=(V,E)$ with $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}$ and $(x_{1},x_{2})\in E$ if and only if $p(x_{1},x_{2})>0$. By (IV.C.19), for each $(x_{1},x_{2})\in E$, we color it by the unique $x_{3}\in\mathcal{X}_{3}$. Assume $|\mathcal{X}_{3}|=k$ and $G$ has $t$ connected components. We denote the number of the vertices in $\mathcal{X}_{i}$ by $n_{i}^{(j)},i=1,2,j=1,2,\cdots,t$ in the connected compoent $C_{j}$ and the probability mass of $C_{j}$ by $p_{j}$. In light of (IV.C.20) and (IV.C.21), equating (IV.C.5) and (IV.C.6), we have
@@
$a\geq\sum\limits_{i=1}^{t}p_{i}\log{k}=\log{k}.$ (IV.C.37)
Note that $H(X_{3})\leq\log{k}$. Thus, by (IV.C.33) and (IV.C.37),
$\log{k}\geq a\geq\log{k},$ (IV.C.38)
which implies that $a$ can only take $\log{k}$ for some positive integer $k$.
The “if ” part of the theorem is immediately implied by Lemma 2 and the fact that the $a=\log{k}$ in $U_{2,3}^{123}$ and whole ray $U_{1,3}^{124}$ are entropic. ∎
@@
For $F=(U_{2,3}^{123},U_{2,3}^{124})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log{k_{1}}$ and $b=\log{k_{2}}$ for positive integers $k_{1},k_{2}$.
$a$$b$Olog2log3log4log2log3log4
@@
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
@@
As (IV.C.66) holds for any $x_{3}^{\prime}\in\mathcal{X}_{3}$, $X_{3}$ is uniformly distributed on $\mathcal{X}_{3}$, and so $H(X_{3})=\log{k_{1}}$ where $k_{1}=|\mathcal{X}_{3}|$. By the same argument, $X_{4}$ is uniform, and so $H(X_{4})=\log{k_{2}}$ where $k_{2}=|\mathcal{X}_{4}|$.
For the “if” part of the theorem, as $\mathbf{h}\in U_{2,3}^{123}$ and $\mathbf{h^{\prime}}\in U_{2,3}^{124}$ are entropic if $a=\log{k_{1}},b=\log{k_{2}}$ for integers $k_{1},k_{2}>0$ [29]. By Lemma 2, all $(\log{k_{1}},\log{k_{2}})$ are entropic. ∎
@@
For $F=(U_{2,3}^{123},U_{1,4})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a\geq 0,b>0$ or $(a,b)=(\log{k},0)$ for positive integer $k$.
$a$$b$Olog2log3log4
@@
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),K\subseteq\{1,2,3\},\mid K\mid=2,\{i,j\}=N_{4}\setminus K.$
@@
Consider the tripartite graph $G=(V,E)$ with $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}\cup\mathcal{X}_{3}$ and $(x_{i},x_{j})\in E$ if and only if $p(x_{i},x_{j})>0$, ${i,j}\in N_{3},i\neq j$. By (IV.C.84)-(IV.C.86), for each $(x_{i},x_{j})\in E$, we color it by the unique $x_{4}\in\mathcal{X}_{4}$. Assume $G$ has $t$ connected components. We denote the number of the vertices in $\mathcal{X}_{i}$ by $n_{i}^{(j)},i=1,2,3,j=1,2,\cdots,t$ in the connected component $C_{j}$ and the probability mass of $C_{j}$ by $p_{j}$. Now for any $(x_{1},x_{2})\in\mathcal{X}_{1}\times\mathcal{X}_{2}$ with $p(x_{1},x_{2})>0$, by (IV.C.84), there exist unique $x_{3}$ and $x_{4}$ with $p(x_{1},x_{2},x_{3},x_{4})>0$, and so by (IV.C.94), $p(x_{1})=p(x_{2})$. Similarly, for $p(x_{1},x_{3})>0$ and $p(x_{2},x_{3})>0$, we have $p(x_{1})=p(x_{3})$ and $p(x_{2})=p(x_{3})$. Then we conclude that $p(x_{i})$=$p(x_{j})$ when $p(x_{i},x_{j})>0$. Thus each vertex in a component of $G$ has the same probability mass. By (IV.C.88)-(IV.C.90), all the edges adjacent to a vertex have the same color, which implies edges in the whole component have the same color. Hence
@@
Now we prove that for all $\mathbf{h}=(a,b)\in F$, $a\geq 0,b>0$ is entropic. Consider $a=\log{k}$ for some positive integer $k$ and $b>0$. Let $X_{4}$ be any random variable on $\mathcal{X}_{4}=\{1,2,\ldots,t\}$ with $H(X_{4})=b$. Note that its probability mass is $p_{i},i\in\mathcal{X}_{4}$. Thus
$b=\sum_{i=1}^{t}-p_{i}\log p_{i}.$ (IV.C.106)
Now assume in each connected component $C_{j}$ of $G$, $\mathcal{X}_{i}^{(j)}$ has $k$ element, that is, $n^{(j)}=k$ for all $j\in\mathcal{X}_{4}$. Let $X^{(j)}_{1}$ and $X^{(j)}_{2}$ be independent and uniformly distributed on $\{0,1,\dots,k-1\}$, and $X^{(j)}_{3}=X^{(j)}_{1}+X^{(j)}_{2}\mod k$. Let $X_{i}=(X^{(j)}_{i},j\in\mathcal{X}_{4})$, $i=1,2,3$. Then
$H(X_{1})$ $=H(X_{2})=H(X_{3})=\sum_{j=1}^{m}-p_{j}\log\frac{p_{j}}{k}$ (IV.C.107)
$=\sum\limits_{j=1}^{m}p_{j}\log{k}+\sum_{i=1}^{m}-p_{j}\log p_{j}$ (IV.C.108)
$=\log{k}+H(X_{4})=a+b.$ (IV.C.109)
@@
Consider $a\neq\log{k},b>0$. Let $c=\lfloor 2^{a}\rfloor$. Let $X_{4}$ be a random variable on $\mathcal{X}_{4}=\{1,2,\ldots,t\}$ with $H(X_{4})=b$. Let $n^{(j)}=c$ for $j\in\mathcal{X}_{4}-t$ and $n^{(t)}=c+l$ for some sufficiently large $l$. Let $X^{(j)}_{1}$ and $X^{(j)}_{2}$ be independent and uniformly distributed on $\{0,1,\dots,c-1\}$ for $j\in\mathcal{X}_{4}-t$ and $X^{(j)}_{3}=X^{(j)}_{1}+X^{(j)}_{2}\mod c$. Let $X^{(t)}_{1}$ and $X^{(t)}_{2}$ be independent and uniformly distributed on $\{0,1,\dots,c+l-1\}$ and $X^{(t)}_{3}=X^{(t)}_{1}+X^{(t)}_{2}\mod c+l$. Let $X_{i}=(X^{(j)}_{i},j\in\mathcal{X}_{4})$, $i=1,2,3$. Then,
@@
$=\sum\limits_{i=1}^{t-1}(-p_{i}\log(\dfrac{p_{i}}{c}))-p_{t}\log\dfrac{p_{t}}{c+l}$ (IV.C.111)
$=\sum\limits_{i=1}^{t}-p_{i}\log(p_{i})+\sum\limits_{i=1}^{t-1}p_{i}\log c+p_{t}\log(c+l)$ (IV.C.112)
$=H(X_{4})+\log{c}({1-p_{t}})+\log{(c+l)}p_{t}.$ (IV.C.113)
@@
Hence, all $\mathbf{h}=(a,b)\in F$ are entropic when $a\geq 0,b>0$. Note that $\mathbf{h}\in U_{2,3}^{123}$ are entropic if and only if $a=\log{k}$ for positive integer $k$. The theorem is proved. ∎

### v3

@@ IV-B Entropy functions on the faces extended from $2$-dimensional faces of $\Gamma_{3}$
By restriction, the above $3$ faces of $\Gamma_{4}$ are mapped to a subset of $\Gamma_{3}$ on $\{2,3,4\}$. The face $(\hat{U}_{2,5}^{1},U_{1,1}^{1})$ of $\Gamma_{4}$ is mapped to the extreme ray $U_{2,3}$ of $\Gamma_{3}$ on $\{2,3,4\}$, which implies that $a$ can only take $\log v$ for integer $v>0$. For the faces $(\hat{U}_{2,5}^{1},U_{1,1}^{2})$ and $(\hat{U}_{2,5}^{1},U_{1,2}^{12})$, by restriction, they are both mapped to the face $(U_{2,3},U^{1}_{1,1})$ of $\Gamma_{3}$ on $\{2,3,4\}$, which implies that $a$ can only take $\log v$ for positive integer $v$ as well.
For the “if” part, as the whole ray $U_{1,1}^{i}$ and $U_{1,2}^{12}$ are entropic, by Lemma 2 and Theorem II-C.4, all $a=\log v$ are entropic. ∎
@@
Let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are the rank functions of the minimal integer polymatroids on the two extreme rays of the face, respectively. Restricting $\mathbf{h}$ on $\{2,3,4\}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}_{1}+b\mathbf{r}^{\prime}_{2}$, where $\mathbf{r}^{\prime}_{i},i=1,2$ are the restriction of $\mathbf{r}_{i}$ on $\{2,3,4\}$. By restriction, The face $(\hat{U}_{2,5}^{1},U_{1,3}^{123})$ of $\Gamma_{4}$ is mapped to the face $(U_{2,3},U^{12}_{1,2})$ of $\Gamma_{3}$ on $\{2,3,4\}$, Thus, $\mathbf{h}^{\prime}\in(U_{2,3},U^{12}_{1,2})$ and $a+b\geq\log\lceil 2^{a}\rceil$ by [10, Theorem 1].
For the “if ” part , we only need to show that $\mathbf{h}$ with $a+b=\log v$ is entropic for each positive integer $v$, then the theorem is implied by Lemma 2 and the fact that $b\mathbf{r}_{2}$ is entropic for all $b\geq 0$. Let $X_{2}$ be uniformly distributed on $\mathbb{I}_{v}$. Let $X_{2}$ and $X_{4}$ be independent, and $X_{4}$ be distributed on $\mathbb{I}_{v}$ such that $H(X_{4})=a$. Let $X_{1}=vX_{2}+X_{4}$, $X_{3}=X_{2}+X_{4}\mod v$. It can be checked that the entropy function of such constructed $(X_{i},i\in N_{4})$ is $(a,\log v-a)$. ∎
@@
For $F=(\hat{U}_{3,5}^{1},U_{1,1}^{1})$, $(\hat{U}_{3,5}^{1},U_{1,1}^{2})$, or $(\hat{U}_{3,5}^{1},U_{1,2}^{12})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log v$ for some integer $v>0$.
@@
For $x_{134}$ with $p(x_{134})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})=p(x_{1},x_{4})$
@@
which implies $p(x_{3})=p(x_{4})$. Since $X_{3}$ and $X_{4}$ are independent, by Lemma 1, they are uniformly distributed on $\mathcal{X}_{3}$ and $\mathcal{X}_{4}$, respectively, and $H(X_{3})=H(X_{4})=\log v$, where $v=|\mathcal{X}_{3}|=|\mathcal{X}_{4}|$. As $(X_{i},i\in\{1,3,4\})$ is its characterizing random vector, we have
@@
Thus $a$ can only take $\log v$.
For the “if” part, it can be proven by Lemma 2 and the fact that $a=\log{v}$ in $\hat{U}_{3,5}^{1}$ and whole ray $U^{1}_{1,1}$, $U^{2}_{1,1}$ and $U_{1,2}^{12}$ are entropic. ∎
@@ IV-C Entropy functions on the faces involving $U_{2,3}^{123}$
For $F=(U_{2,3}^{123},U_{1,3}^{124})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v>0$.
@@
$H(X_{3j})$ $=H(X_{3})+H(X_{j}),j\in N_{4}\setminus\{3\},$
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),k\neq 3,$
$\qquad\qquad\qquad\qquad\{i,j,k\}\neq\{1,2,3\},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),K\subseteq\{1,2,3\}$
$\qquad\qquad\quad|K|=2,\{i,j\}=N_{4}\setminus K.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
@@
Consider the bipartite graph $G=(V,E)$ with $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}$ and $(x_{1},x_{2})\in E$ if and only if $p(x_{1},x_{2})>0$. By (IV.C.19), for each $(x_{1},x_{2})\in E$, we color it by the unique $x_{3}\in\mathcal{X}_{3}$. Assume $|\mathcal{X}_{3}|=v$ and $G$ has $t$ connected components. We denote the number of the vertices of $\mathcal{X}_{i}$ in the connected component $C_{j}$ by $n_{i}^{(j)},i=1,2,j=1,2,\cdots,t$ and the probability mass of $C_{j}$ by $p_{j}$, that is, the probability of the event that the random vector takes a tuple in $C_{j}$. In light of (IV.C.20) and (IV.C.21), equating (IV.C.5) and (IV.C.6), we have
@@
$a\geq\sum\limits_{i=1}^{t}p_{i}\log{v}=\log{v}.$ (IV.C.37)
Note that $H(X_{3})\leq\log{v}$. Thus, by (IV.C.33) and (IV.C.37),
$\log{v}\geq a\geq\log{v},$ (IV.C.38)
which implies that $a$ can only take $\log{v}$ for some positive integer $k$.
The “if” part of the theorem is immediately implied by Lemma 2 and the fact that the $a=\log{v}$ in $U_{2,3}^{123}$ and whole ray $U_{1,3}^{124}$ are entropic. ∎
@@
For $F=(U_{2,3}^{123},U_{2,3}^{124})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log{v_{1}}$ and $b=\log{v_{2}}$ for positive integers $v_{1},v_{2}$.
@@
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
@@
As (IV.C.66) holds for any $x_{3}^{\prime}\in\mathcal{X}_{3}$, $X_{3}$ is uniformly distributed on $\mathcal{X}_{3}$, and so $H(X_{3})=\log{v_{1}}$ where $v_{1}=|\mathcal{X}_{3}|$. By the same argument, $X_{4}$ is uniform, and so $H(X_{4})=\log{v_{2}}$ where $v_{2}=|\mathcal{X}_{4}|$.
For the “if” part of the theorem, as $\mathbf{h}\in U_{2,3}^{123}$ and $\mathbf{h^{\prime}}\in U_{2,3}^{124}$ are entropic if $a=\log{v_{1}},b=\log{v_{2}}$ for integers $v_{1},v_{2}>0$ [35]. By Lemma 2, all $(\log{v_{1}},\log{v_{2}})$ are entropic. ∎
@@
For $F=(U_{2,3}^{123},U_{1,4})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a\geq 0,b>0$ or $(a,b)=(\log{v},0)$ for positive integer $v$.
[…]
@@
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),K\subseteq\{1,2,3\},$
$\qquad\qquad\quad|K|=2,\{i,j\}=N_{4}\setminus K.$
@@
Consider the tripartite graph $G=(V,E)$ with $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}\cup\mathcal{X}_{3}$ and $(x_{i},x_{j})\in E$ if and only if $p(x_{i},x_{j})>0$, ${i,j}\in N_{3},i\neq j$. By (IV.C.84)-(IV.C.86), for each $(x_{i},x_{j})\in E$, we color it by the unique $x_{4}\in\mathcal{X}_{4}$. Assume $G$ has $t$ connected components. We denote the number of the vertices of $\mathcal{X}_{i}$ in the connected component $C_{j}$ by $n_{i}^{(j)},i=1,2,3,j=1,2,\cdots,t$ and the probability mass of $C_{j}$ by $p_{j}$. Now for any $(x_{1},x_{2})\in\mathcal{X}_{1}\times\mathcal{X}_{2}$ with $p(x_{1},x_{2})>0$, by (IV.C.84), there exist unique $x_{3}$ and $x_{4}$ with $p(x_{1},x_{2},x_{3},x_{4})>0$, and so by (IV.C.94), $p(x_{1})=p(x_{2})$. Similarly, for $p(x_{1},x_{3})>0$ and $p(x_{2},x_{3})>0$, we have $p(x_{1})=p(x_{3})$ and $p(x_{2})=p(x_{3})$. Then we conclude that $p(x_{i})$$p(x_{j})$ when $p(x_{i},x_{j})>0$. Thus each vertex in a component of $G$ has the same probability mass. By (IV.C.88)-(IV.C.90), all the edges adjacent to a vertex have the same color, which implies edges in the whole component have the same color. Hence
@@
Now we prove that for all $\mathbf{h}=(a,b)\in F$, $a\geq 0,b>0$ is entropic. For $a=\log v$ with some positive $v$ and $b>0$, the fact that $(a,b)$ is entropic can be implied by Lemma2. Here we give a constructive proof of the case. Let $X_{4}$ be any random variable on $\mathcal{X}_{4}=\{0,1,\ldots,t-1\}$ with $H(X_{4})=b$. Note that its probability mass is $p_{i},i\in\mathcal{X}_{4}$. Thus
$b=\sum_{i=}^{t-1}-p_{i}\log p_{i}.$ (IV.C.106)
Now assume in each connected component $C_{j}$ of $G$, $\mathcal{X}_{i}^{(j)}$ has $k$ element, that is, $n^{(j)}=k$ for all $j\in\mathcal{X}_{4}$. Let $X^{(j)}_{1}$ and $X^{(j)}_{2}$ be independent and uniformly distributed on $\mathbb{I}_{v}$, and $X^{(j)}_{3}=X^{(j)}_{1}+X^{(j)}_{2}\mod v$. Let $X_{i}=(X^{(j)}_{i},j\in\mathcal{X}_{4})$, $i=1,2,3$. Then
$H(X_{1})$ $=H(X_{2})=H(X_{3})=\sum_{j=0}^{t-1}-p_{j}\log\frac{p_{j}}{k}$ (IV.C.107)
$=\sum\limits_{j=0}^{t-1}p_{j}\log{v}+\sum_{j=0}^{t-1}-p_{j}\log p_{j}$ (IV.C.108)
$=\log{v}+H(X_{4})=a+b.$ (IV.C.109)
@@
Consider $a\neq\log{v},b>0$. Let $c=\lfloor 2^{a}\rfloor$. Let $X_{4}$ be a random variable on $\mathcal{X}_{4}=\{0,1,\ldots,t-1\}$ with $H(X_{4})=b$. Let $n^{(j)}=c$ for $j\in\mathcal{X}_{4}\setminus\{t-1\}$ and $n^{(t-1)}=c+l$ for some sufficiently large $l$. Let $X^{(j)}_{1}$ and $X^{(j)}_{2}$ be independent and uniformly distributed on $\{0,1,\dots,c-1\}$ for $j\in\mathcal{X}_{4}\setminus\{t-1\}$ and $X^{(j)}_{3}=X^{(j)}_{1}+X^{(j)}_{2}\mod c$. Let $X^{(t-1)}_{1}$ and $X^{(t-1)}_{2}$ be independent and uniformly distributed on $\{0,1,\dots,c+l-1\}$ and $X^{(t-1)}_{3}=X^{(t-1)}_{1}+X^{(t-1)}_{2}\mod c+l$. Let $X_{i}=(X^{(j)}_{i},j\in\mathcal{X}_{4})$, $i=1,2,3$. Then,
@@
$=\sum\limits_{i=0}^{t-2}(-p_{i}\log(\dfrac{p_{i}}{c}))-p_{t-1}\log\dfrac{p_{t-1}}{c+l}$ (IV.C.111)
$=\sum\limits_{i=0}^{t-1}-p_{i}\log(p_{i})+\sum\limits_{i=0}^{t-2}p_{i}\log c+p_{t-1}\log(c+l)$ (IV.C.112)
$=H(X_{4})+\log{c}({1-p_{t-1}})+\log{(c+l)}p_{t-1}.$ (IV.C.113)
@@
Hence, all $\mathbf{h}=(a,b)\in F$ are entropic when $a\geq 0,b>0$. Note that $\mathbf{h}\in U_{2,3}^{123}$ are entropic if and only if $a=\log{v}$ for positive integer $v$. The theorem is proved. ∎
Theorem IV-C.3 can be considered as a corollary of [32, Lemma 4], which is restated as [36, Lemma 6].